# Learning incomplete factorization preconditioners for GMRES

### v1

@@ 3.2 Model training
The goal of the learned preconditioner ${\bm{P}}={\bm{L}}{\bm{U}}$ is to improve the spectral properties of the preconditioned linear equation system ${\bm{A}}{\bm{P}}^{-1}$. However, directly optimizing the condition number $\kappa$ of the system as suggested by sappl2019deep is computationally infeasible for large-scale problems. For computational tractability, we are also not taking the whole spectrum of the equation system into account but concentrate on the edges of the spectrum. In other words, we consider only the largest and smallest singular values of the preconditioned equation system.
@@
The largest singular value of the matrix ${\bm{P}}^{-1}{\bm{A}}$ is upper bounded by the Frobenius norm: $\sigma_{\max}({\bm{A}}{\bm{P}}^{-1})\leq\epsilon^{-1}\;\|{\bm{A}}-{\bm{P}}\|_{F}+1$.
@@
where ${\bm{w}}$ is a standard normal distributed vector. This loss function was first suggested by hausner2023neural as an approximation of the Froebnius norm.
@@
The drawback of this loss is that it requires computing ${\bm{A}}^{-1}{\bm{w}}$ during training. In other words, we need to solve many potentially ill-conditioned linear systems in order to train the learned preconditioner. chen2024graph proposes a similar loss function and approximates the solutions to the linear equation system instead using Arnoldi’s method.
To maintain computational efficiency we can make use of a supervised dataset instead, consisting of tuples $({\bm{A}}_{i},{\bm{x}}_{i},{\bm{b}}_{i})$ such that ${\bm{A}}_{i}{\bm{x}}_{i}={\bm{b}}_{i}$. This requires us to solve each training problem only once which can be computed beforehand in an offline fashion. We replace the standard distributed samples ${\bm{w}}$ in Equation (6) with the available training samples. This avoids solving problem instances during training but leads to a biased approximation of the previous loss function:
@@
li2023learning originally suggested this loss function as an approximation of the Frobenius norm loss and motivate it by including inductive bias from the training data. Later, trifonov2024learning obtain the same approximation by reweighting of Equation (5) with the inverse matrix. They conjecture the connection of the loss function with small singular values and, further, empirically show that minimizing this loss increases the smallest singular value of the preconditioned system.
By combining the two previous loss functions from Equation (6) and the unregularized version of Equation (5), we introduce a novel combined loss with the goal to further improve the conditioning by taking into account both ends of the spectrum. The combined loss function and the approximation using a supervised dataset similar to Equation (7) are
@@
where $\alpha$ is a hyper-parameter chosen to balance the two individual loss terms. During the training, we are minimizing the empirical risk over the corresponding unsupervised or supervised dataset to learn the parameters:
$\hat{{\bm{\theta}}}\in\operatorname*{arg\,min}_{\bm{\theta}}\sum_{i=1}^{|D|}\mathcal{L}({\bm{P}}_{\bm{\theta}};{\bm{A}}_{i},{\bm{x}}_{i},{\bm{b}}_{i}).$ (9)
The proofs for the lemmas and the connection of the loss functions to previous methods from the literature can be found in Appendix C.
Table 1: Average performance of classical and data-driven preconditioners with different loss functions evaluated on $10$ test samples. We show the conditioning of the right-preconditioned system ${\bm{A}}{\bm{P}}^{-1}$ as well as the convergence of the GMRES algorithm in terms of the total time required to compute the preconditioner and solve the systems and the number of iterations. For the unpreconditioned method, we compute the loss for ${\bm{P}}={\bm{I}}$. Arrows indicate if higher or lower is better. The best performance is marked in bold.
@@ 4 Experiments & Results
Comparison of loss functions
Figure 3: Distribution of singular values of the preconditioned system for different choices of preconditioner.
In Table 1, we present the performance of the learned preconditioner given different loss functions corresponding to the bounds derived in the previous section. We can see that minimizing $\mathcal{L}_{\max}$ and $\mathcal{L}_{\min}$ decrease the largest and increase the smallest singular value of the system respectively and the models achieve the lowest loss for the respective functions. This shows that during training, we are able to minimize the loss functions and by that implicitly optimize the bounds for the singular values.
Optimizing the biased approximation ^$\hat{\mathcal{L}}_{\min}$ still performs well in terms of decreasing the original loss function corresponding to the bound of the smallest singular value. The resulting preconditioning matrix is highly ill-conditioned for the chosen $\epsilon$ parameter leading to numerical instabilities in the inverseion of ${\bm{P}}$ and inaccurate results for the singular value decomposition due to the overfitting of the model on the training data.
@@
Note that, the problem at hand is unbalanced as the spectrum of the original system is already skewed towards small singular values. The performance in terms of GMRES iterations of the different loss functions also depends on the initial distribution of singular values. In Figure 3 we show this distribution for different preconditioners compared to the original distribution.
Comparison with classical preconditioners  Additionally, we compare in Table 1 the performance of the learned preconditioner with the classical Jacobi and incomplete LU method without fill-ins, as well as, the case where no preconditioner is applied. In general, the results show the vital importance of selecting an appropriate preconditioner. For the problem considered, ILU and the learned preconditioner perform nearly on par. We can see that even though the conditioning in terms of the largest and smallest singular value of the ILU method is by far inferior to the other preconditioners, it still manages to substantially reduce the number of iterations for convergence.
This shows that the spectrum edges, while important, are not the only factor affecting the method’s efficiency. It underscores the importance of considering the entire spectrum of the preconditioned system when analyzing the method’s convergence.
@@ 5 Discussion & Conclusion
In this paper, we learn incomplete factorizations of sparse matrices directly from data. When applied as a preconditioner our approach leads to a significant reduction in terms of iterations compared to the unpreconditioned GMRES. Combining data-driven methods with existing algorithms has a huge potential since it allows obtaining the best results from both worlds: guarantees about the convergence is obtained from classical analysis of the algorithms while the data-driven components allow us to learn directly from data to improve upon the existing algorithms.
Limitations  Our approach assumes having access to a distribution of linear equation systems that share similarities between the problem instances. However, in practice, it is not always trivial to obtain such a distribution of problems. Further, the initial cost of training the network needs to be amortized over many future problems that need to be solved. However, given the initial sample problems, new problems can be solved faster in an online fashion which is important for time-critical applications.
Future work  The GNN is designed so that the learned preconditioner matches the sparsity pattern of the input matrix ${\bm{A}}$. While dynamically learning the sparsity pattern and matrix elements could improve preconditioner quality, the large search space and combinatorial complexity make this challenging.
Our proposed loss function accounts for the spectrum edges in the preconditioned system. However, beyond the edges and condition number, the distribution of singular values also affects the convergence of Krylov subspace methods. Designing loss functions that consider this can further improve the performance of learned preconditioners.
@@ Acknowledgments
We would like to thank Daniel Hernández Escobar and Sebastian Mair for helpful discussions and feedback. This work was supported by the the Göran Gustafsson Foundation and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix A GMRES algorithm
The right-preconditioned GMRES method is shown in Algorithm 1. The non-preconditioned version can be obtained by using the identity matrix as a preconditioner (${\bm{P}}={\bm{I}}$). Operations where the preconditioner is applied are highlighted for better readability in this section.
@@
where ${\bm{V}}_{k}$ consists of the first $k$ columns of ${\bm{V}}_{k+1}$, forming an orthonormal basis for the Krylov subspace.
Based on the factorization of this basis, the residual and approximate solution can be efficiently obtained by solving a smaller and upper-triangular system derived from the QR factorization of ${\bm{H}}_{k,k+1}$, as shown in line 20 of the GMRES algorithm.
@@
In the case of preconditioner given by triangular factorizations such as ILU and our learned approach, the matrix is obtained by solving the system ${\bm{L}}{\bm{U}}{\bm{v}}={\bm{r}}$ which can be achieved in $\mathcal{O}(n^{2})$ operations using the forward-backward substitution and can be further accelerated by exploiting the sparse structure of the matrices. The solution to solving these systems is equivalent to explicitly inverting the matrices and computing the matrix-vector product ${\bm{U}}^{-1}{\bm{L}}^{-1}{\bm{r}}$ which forms the preconditioning matrix as ${\bm{P}}^{-1}=({\bm{L}}{\bm{U}})^{-1}$. However, solving the triangular systems is more efficient in practice.
@@ B.1 Dataset
Our goal with the provided dataset is not solving a real-world problem. Rather, we focus on a highly ill-conditioned synthetic dataset which allows us to systematically test the different loss functions derived. For the problem scale used, direct methods are much superior compared to the shown results.
@@
$\begin{split}-\nabla^{2}u(x)&=f(x)\quad x\in\Omega,\\ u(x)&=u_{D}(x)\quad x\in\partial\Omega,\end{split}$ (11)
@@
Noisy data  However, we are not specifically interested in spd matrices. Therefore, to obtain a distribution $\mathbb{A}$ over a large number of Poisson equation PDE problems that can be efficiently solved using GMRES, we perturb all the nonzero entries with standard Gaussian noise, which yields arbitrary matrices not necessarily spd. Let us denote the perturbed matrix as ${\bm{A}}^{\prime}=(a^{\prime}_{ij})$.
We can represent this mathematically by:
@@ B.2 Network architecture
Inputs and transformations  As described in Section 3, we transform the original linear equation system matrix ${\bm{A}}$ into a graph $\mathcal{G}$ using the Coates graph representation. We use the nonzero elements of the matrix as edge features and add a second edge feature as a positional encoding of the output. For the node features, we utilize the same set of $8$ input features as previously applied by hausner2023neural describing the structural properties of the matrix.
@@
Output  The final edge embedding is chosen to be scalar by designing the edge update function to produce a single output, i.e. $e_{ij}^{(L)}\in\mathbb{R}$, allowing us to directly transform the output into a lower- and upper-triangular matrix. We enforce the diagonal elements of ${\bm{U}}$ to be ones and apply the activation function $\phi$ from Equation (3) to the diagonal elements of ${\bm{L}}$. Note that during training, we instead use the approximation $\hat{\phi}$ from Equation (4) in line 25 as discussed in the main text.
@@
21:  $\triangleright$ Form the lower triangular matrix $\bm{L}=(l_{ij})$
$\left(l_{ij}\right)\leftarrow\left\{\begin{array}[]{ll}e_{ij}^{(L)},&\text{if }(i,j)\in\mathcal{E}\text{ and }i>j,\\ 0,&\text{otherwise}\end{array}\right.$
22:  $\triangleright$ Form the upper triangular matrix $\bm{U}=(u_{ij})$
$\left(u_{ij}\right)\leftarrow\left\{\begin{array}[]{ll}e_{ij}^{(L)},&\text{if }(i,j)\in\mathcal{E}\text{ and }i<j,\\ 0,&\text{otherwise}\end{array}\right.$
@@
For loss functions that require solving the system ${\bm{A}}^{-1}{\bm{w}}$, training takes significantly longer with approximately $20$ minutes. However, solving the systems can be implemented efficiently for the problem scale that we consider in our experiments.
Tuning  We only conducted a minimal level of hyper-parameter tuning for the model architecture. Tuning the loss function in Equation (8), optimization of the objective becomes numerically more difficult as the two terms are in conflict with each other. We choose the hyper-parameter $\alpha=\frac{1}{5}$ based on the validation set performance to balance out the two loss terms of the combined loss.

### v2

@@ 3.2 Model training
The goal of the learned preconditioner ${\bm{P}}={\bm{L}}{\bm{U}}$ is to improve the spectral properties of the preconditioned linear equation system ${\bm{A}}{\bm{P}}^{-1}$. However, as sappl2019deep suggest, directly optimizing the system’s condition number $\kappa$ becomes computationally intractable for large-scale problems. For computational tractability, we are also not taking the whole spectrum of the equation system into account but concentrating on the edges of the spectrum. In other words, we consider only the largest and smallest singular values of the preconditioned equation system.
@@
The largest singular value of the matrix ${\bm{A}}{\bm{P}}^{-1}$ is upper bounded by the Frobenius norm: $\sigma_{\max}({\bm{A}}{\bm{P}}^{-1})\leq\epsilon^{-1}\;\|{\bm{A}}-{\bm{P}}\|_{F}+1$.
@@
where ${\bm{w}}$ is a standard normal distributed vector. This loss function has been previously applied by hausner2023neural as an unbiased estimator of the squared Frobenius norm distance.
@@
The drawback of this loss is that it requires computing ${\bm{A}}^{-1}{\bm{w}}$ during training. In other words, training the learned preconditioner requires solving many potentially ill-conditioned linear systems.
To maintain computational efficiency and avoid the need to solve equation systems online during training, we can instead utilize a supervised dataset composed of tuples $({\bm{A}}_{i},{\bm{x}}_{i},{\bm{b}}_{i})$, where each tuple satisfies ${\bm{A}}_{i}{\bm{x}}_{i}={\bm{b}}_{i}$. This requires us to solve each training problem only once which can be computed beforehand in an offline fashion. We replace the standard distributed samples ${\bm{w}}$ in Equation (6) with the available training samples, resulting in a biased approximation of the original loss function:
@@
This loss function has been previously introduced by li2023learning as an inductive bias introduced by the dataset for the Frobenius norm. Recently, trifonov2024learning derived this loss function and the unbiased variant from Equation (6) based on reweighing the result from Lemma 1 using the inverse matrix as weights. They conjecture that optimizing this loss leads to minimizing the low-frequency components in the preconditioned equation system resulting in larger small singular values.
By combining the two previous loss functions from Equation (6) and Equation (5), we introduce a novel combined loss with the goal to further improve the conditioning by taking into account both ends of the spectrum. The combined loss function, along with its approximation using a supervised dataset similar to Equation (7), is defined as:
@@
where $\alpha$ is a hyper-parameter that controls the trade-off between the two loss components. During training, we are minimizing the empirical risk over the corresponding unsupervised or supervised dataset to learn the parameters:
$\hat{{\bm{\theta}}}\in\operatorname*{arg\,min}_{\bm{\theta}}\sum_{i=1}^{|\mathcal{D}|}\mathcal{L}({\bm{P}}_{\bm{\theta}};{\bm{A}}_{i},{\bm{x}}_{i},{\bm{b}}_{i}).$ (9)
The proofs for the lemmas and the Froebnius norm approximation can be found in Appendix C.
Table 1: Average performance of classical and data-driven preconditioners with different loss functions evaluated on $10$ test samples. We show the conditioning of the right-preconditioned system ${\bm{A}}{\bm{P}}^{-1}$ in the first three columns indicating the smallest ($\sigma_{\min}$) and largest ($\sigma_{\max}$) singular values as well as the condition number ($\kappa$). For each preconditioner, we compute different Frobenius norm loss functions proportional to the bounds in Lemma 1 and Lemma 2. Further, we compare the convergence of the preconditioner in the GMRES algorithm in terms of the total time (including both the time to compute the preconditioner and solve the systems) and the number of iterations. For the unpreconditioned method, we compute the loss for ${\bm{P}}={\bm{I}}$.
@@ 4 Experiments & Results
Comparison of loss functions  In Table 1, we present the performance of the learned preconditioner given different loss functions corresponding to the bounds derived in the previous section. We can see that minimizing $\mathcal{L}_{\max}$ and $\mathcal{L}_{\min}$ decrease the largest and increase the smallest singular value of the system respectively and the models achieve the lowest loss for the respective functions. This shows that during training, we are able to minimize the loss functions and by that implicitly optimize the bounds for the singular values.
Optimizing the biased approximation ^$\hat{\mathcal{L}}_{\min}$ still performs well in terms of decreasing the original unbiased loss function. However, when computing the preconditioner inverse and singular value decomposition, numerical instabilities prevent the application of the learned output due to poor conditioning. This leads, in turn, to significantly worse performance in terms of GMRES iterations.
@@
Note that, the problem at hand is unbalanced as the spectrum of the original system is already skewed towards small singular values. In terms of GMRES iterations, the performance of the different loss functions also depends on the initial distribution of singular values. In Figure 3 we show this distribution for different preconditioners compared to the original distribution. Note that the original distribution is shifted to small singular values. We discuss the results in more detail in Appendix D.1.
Figure 3: Descending log-scale plot of ordered singular values of a single problem instance for learned preconditioners derived from the different loss functions.
Comparison with data-driven preconditioners  We additionally compare our method with previously developed data-driven preconditioners for incomplete Cholesky factorizations (Learned IC) with an $\exp$-activation function for the diagonal elements to ensure invertibility. We use the loss function $\mathcal{L}_{\text{max}}$ to train the method, similar to previous approaches (hausner2023neural), and choose the same hyper-parameters as our model for a fair comparison. The details of the baseline architecture are provided in Appendix B.3.
The results indicate that, although the Learned IC method reduces the number of iterations compared to the unpreconditioned GMRES solver, it performs worse in terms of both loss and conditioning when compared to our proposed method. Further, the GMRES performance is significantly worse than our proposed LU factorization-based techniques. This is, however, not very surprising as the LU factorization offers more flexibility to approximate the true matrix factorization as two separate factors are learned.
Comparison with classical preconditioners  Additionally, we compare in Table 1 the performance of the learned preconditioner against the classical Jacobi and incomplete LU method without fill-ins, as well as the case where no preconditioner is applied. In general, the results show the vital importance of selecting an appropriate preconditioner. For the problem considered, ILU and the learned preconditioner perform nearly on par in terms of required iterations to solve the problem instances. We can see that even though the conditioning in terms of the largest and smallest singular value of the ILU method is by far inferior to the other preconditioners, it still manages to substantially reduce the number of iterations for convergence.
This shows that the spectrum edges, while important, are not the only factor affecting the method’s efficiency. Further, this underscores the importance of considering the entire spectrum of the preconditioned system when analyzing the method’s convergence as well as the alignment of the right-hand side ${\bm{b}}$ of the system with the singular vectors of ${\bm{A}}$ (benzi2002preconditioning; carson2024towards).
@@ 5 Discussion & Conclusion
In this paper, we learn incomplete LU factorizations of sparse matrices directly from data using graph neural networks. When applied as a preconditioner our approach leads to a significant reduction in terms of iterations compared to the unpreconditioned GMRES and performs similar to classical incomplete factorization methods. Combining data-driven methods with existing numerical algorithms has a huge potential since it allows obtaining the best results from both worlds: guarantees about the convergence are obtained from classical analysis of the methods while the data-driven components allow us to learn directly from data to improve upon the existing algorithms on specific data distributions.
Limitations  Our approach assumes having access to a distribution of linear equation systems that share similarities between the problem instances. However, in practice, it is not always trivial to obtain such a distribution of problems. Further, the initial cost of training the network needs to be amortized over many future problems that need to be solved. However, given the initial sample problems, new problems can be solved faster in an online setting which is important for time-critical applications.
In our experiments, we only train and evaluate our method on a single dataset of very limited size which is inspired by numerical computations. Applying the method to more diverse problem classes with different initial distributions of singular values is required in the future to obtain more general results about the performance.
Although the combined loss function $\mathcal{L}_{\text{com}}$ accelerates GMRES convergence, the conflicting loss terms can introduce instability during training. Additionally, finding an appropriate balance between the two components using the hyper-parameter $\alpha$ is challenging, as a poorly chosen value might emphasize one part of the spectrum more than intended, resulting in a similar performance to using one of the loss functions individually.
Future work  Our GNN is designed such that the learned preconditioner matches the sparsity pattern of the input matrix ${\bm{A}}$. While dynamically learning the sparsity pattern and matrix elements could improve preconditioner quality, the large search space and combinatorial complexity make this challenging.
Our proposed loss function accounts for the edges of the spectrum in the preconditioned system. However, beyond the edges and condition number, the distribution of singular values as well as the right-hand side also affects the convergence of Krylov subspace methods (carson2024towards). Designing loss functions that consider these additional aspects can further improve the performance of learned preconditioners.
An interesting idea for future work is to combine our data-driven incomplete LU factorization with classical ILU methods along the lines of the method proposed by trifonov2024learning combining the advantages of classical and data-driven methods. The theoretical analysis of the different loss functions from our results can then be applied to construct novel hybrid preconditioners.
@@ Acknowledgments
We thank Daniel Hernández Escobar and Sebastian Mair for helpful discussions and feedback. This work was supported by the Göran Gustafsson Foundation and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
@@ Appendix A GMRES algorithm
The right-preconditioned GMRES is shown in Algorithm 1. The non-preconditioned version can be obtained by using the identity matrix as a preconditioner (${\bm{P}}={\bm{I}}$). Operations where the preconditioner is applied are highlighted for better readability in this section.
@@
where ${\bm{V}}_{k}$ consists of the first $k$ columns of the matrix ${\bm{V}}_{k+1}$, forming an orthonormal basis for the Krylov subspace.
Based on this factorization of this basis, the residual and approximate solution can be efficiently obtained by solving a smaller and upper-triangular system derived from the QR factorization of ${\bm{H}}_{k,k+1}$, as shown in line 20 of the GMRES algorithm.
@@
In the case of preconditioner given by triangular factorizations such as ILU and our learned approach, the matrix is obtained by solving the system ${\bm{L}}{\bm{U}}{\bm{v}}={\bm{r}}$ which can be achieved in $\mathcal{O}(n^{2})$ operations using the forward-backward substitution and can be further accelerated by exploiting the sparse structure of the matrices. The solution to solving these systems is equivalent to explicitly inverting the matrices and computing the matrix-vector product ${\bm{U}}^{-1}{\bm{L}}^{-1}{\bm{r}}$ which forms the preconditioning matrix as ${\bm{P}}^{-1}=({\bm{L}}{\bm{U}})^{-1}$. However, solving the triangular systems is more efficient in practice as the generally non-sparse inverse matrices do not need to be stored.
Convergence  Obtaining descriptive convergence bound is in general non-trivial due to the complex nature of GMRES. For non-symmetric matrices, singular values provide a more complete understanding of a matrix’s behavior across all directions in space compared to eigenvalues. Eigenvalues are tied to specific directions – eigenvectors – that remain unchanged under transformation by the matrix ${\bm{A}}$. However, non-symmetric matrices may lack sufficient eigenvectors to describe their impact across all directions, particularly if they are defective or not diagonalizable. In contrast, singular values offer a full directional picture. The matrix can be decomposed, using singular value decomposition (SVD), as ${\bm{A}}={\bm{U}}\bm{\Sigma}{\bm{V}}^{\mkern-1.5mu\mathsf{T}}$, where $\bm{\Sigma}$ is a diagonal matrix of singular values, and ${\bm{U}}$ and ${\bm{V}}$ are orthogonal matrices containing left and right singular vectors, respectively. These singular vectors capture ${\bm{A}}$’s scaling effects across all directions, making singular values especially valuable for analyzing convergence and stability in iterative methods that require complete directional information saad86gmres; golub2013matrix. As the generated Krylov subspace in GMRES highly depends on the matrix’s scaling behavior, SVD offers the tools to understand the iterative process through the transformations of ${\bm{A}}$ trefethen1997numerical.
The condition number, $\kappa({\bm{A}})=\frac{\sigma_{\max}({\bm{A}})}{\sigma_{\min}({\bm{A}})}$, defined by the largest and smallest singular values of ${\bm{A}}$ is a popular metric to assess the convergence of iterative methods. This metric describes how well-conditioned the matrix is, with a lower $\kappa({\bm{A}})$ indicating better clustering of singular values, typically around 1 in well-conditioned systems. A smaller $\kappa({\bm{A}})$ often leads to faster convergence since all directions contribute meaningfully to residual reduction. However, if $\kappa({\bm{A}})$ is large due to a very small $\sigma_{\min}({\bm{A}})$, convergence slows as the matrix has limited influence in the directions corresponding to small singular values. This condition can cause stagnation, requiring more iterations to meet convergence criteria as $\kappa({\bm{A}})$ grows saad86gmres; golub2013matrix.
Preconditioning addresses this by adjusting the singular value spectrum, ideally reducing $\kappa({\bm{A}})$ to cluster singular values more effectively and ensuring significant influence in all directions of the subspace. A well-chosen preconditioner modifies ${\bm{A}}$’s singular values to improve conditioning, thus lowering the number of iterations needed to achieve a desired tolerance. This approach makes GMRES particularly effective for systems that are well-conditioned or well-preconditioned, solidifying singular values as the preferred metric over eigenvalues for evaluating convergence and stability in this context benzi2002preconditioning.
Right-hand side  Apart from the condition number and singular value distribution, the right-hand side ${\bm{b}}$ in the system also plays a crucial role in GMRES convergence through its interaction with the singular value decomposition of ${\bm{A}}$ or the preconditioned system. Decomposing ${\bm{b}}=\sum_{i=1}^{n}\alpha_{i}{\bm{u}}_{i}$, with $\{{\bm{u}}_{i}\}$ as the left singular vectors of ${\bm{A}}$, reveals how ${\bm{b}}$ aligns with ${\bm{A}}$’s left-singular values. The coefficients $\alpha_{i}={\bm{u}}_{i}^{\mkern-1.5mu\mathsf{T}}{\bm{b}}$ measure ${\bm{b}}$’s alignment with each left singular vector ${\bm{u}}_{i}$. Strong alignment with directions tied to small singular values $\sigma_{i}$ causes these directions to dominate the residual and slow down GMRES convergence golub2013matrix; benzi2002preconditioning.
Our data-driven preconditioning method indirectly addresses the interaction between the right-hand side ${\bm{b}}$ and the singular values of ${\bm{A}}$ by sampling ${\bm{b}}$ from a distribution in the approximate loss functions. This ensures robustness by forcing the preconditioner to handle diverse right-hand sides and target weakly scaled directions of ${\bm{A}}$. However, our method does not explicitly exploit the alignment of a specific right-hand side ${\bm{b}}$ with the singular vectors of ${\bm{A}}$. To refine this approach, the loss function can be enhanced in future work to incorporate the dependence of a specific right-hand side and its alignment with the singular values explicitly.
@@ B.1 Dataset
Our goal with the provided dataset is not to solve a real-world problem. Rather, we focus on a highly ill-conditioned synthetic dataset which allows us to systematically test the different loss functions derived. For the problem scale used, direct methods are much superior compared to the shown results.
@@
$-\nabla^{2}u(x)$ $=f(x)$ $x\in\Omega,$ (11)
$u(x)$ $=u_{D}(x)$ $x\in\partial\Omega,$ (12)
@@
Noisy data  However, we are not specifically interested in symmetric and positive definite matrices that are generated through finite element discretization. Therefore, to obtain a distribution $\mathbb{A}$ over a large number of Poisson equation PDE problems with more general structure that can be efficiently solved using GMRES, we perturb all the non-zero entries with standard Gaussian noise, which yields arbitrary matrices not necessarily spd. Let us denote the perturbed matrix as ${\bm{A}}^{\prime}=(a^{\prime}_{ij})$. We can represent this mathematically by:
@@ B.2 Network architecture
Inputs and transformations  As described in Section 3, we transform the original linear equation system matrix ${\bm{A}}$ into a graph $\mathcal{G}$ using the Coates graph representation. We use the non-zero elements of the matrix as edge features and add a second edge feature as a positional encoding of the output. For the node features, we utilize the same set of $8$ input features as previously applied by hausner2023neural describing the structural properties of the matrix:
  • the node degree deg(v)
  • the maximum degree of neighboring nodes
  • the minimum degree of neighboring nodes
  • the mean degree of neighboring nodes
  • the variance of degrees of neighboring nodes
  • the diagonal dominance
  • the diagonal decay
  • the position of the node in the linear system
@@
Output  The final edge embedding is chosen to be scalar by designing the edge update function to produce a single output, i.e., $e_{ij}^{(L)}\in\mathbb{R}$, allowing us to directly transform the output into a lower- and upper-triangular matrix. We enforce the diagonal elements of ${\bm{U}}$ to be ones and apply the activation function $\zeta$ from Equation (3) to the diagonal elements of ${\bm{L}}$. Note that during training, we instead use the approximation $\hat{\zeta}_{\epsilon}$ from Equation (4) in line 25 as discussed in the main text. This allows zero diagonal elements but no matrix inversion of ${\bm{P}}$ during training is required.
@@
21:  $\triangleright$ Form the lower triangular matrix $\bm{L}=(l_{ij})$:
$\left(l_{ij}\right)\leftarrow\left\{\begin{array}[]{ll}e_{ij}^{(L)},&\text{if }(i,j)\in\mathcal{E}\text{ and }i>j,\\ 0,&\text{otherwise.}\end{array}\right.$
22:  $\triangleright$ Form the upper triangular matrix $\bm{U}=(u_{ij})$:
$\left(u_{ij}\right)\leftarrow\left\{\begin{array}[]{ll}e_{ij}^{(L)},&\text{if }(i,j)\in\mathcal{E}\text{ and }i<j,\\ 0,&\text{otherwise.}\end{array}\right.$
@@
For loss functions that require solving the system ${\bm{A}}^{-1}{\bm{w}}$, training takes significantly longer with approximately $20$ minutes. However, solving the systems can be implemented efficiently for the problem scale that we consider in our experiments. For real-world examples, the additional time required would increase significantly.
Tuning  We only conducted a minimal level of hyper-parameter tuning for the model architecture. Tuning the loss function in Equation (8), optimization of the objective becomes numerically more difficult as the two terms are in conflict with each other. We choose the hyper-parameter $\alpha=\nicefrac{{1}}{{7}}$ based on the validation set performance to balance out the two loss terms of the combined loss.